OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find [f(4)]². a.81 b.257 c.289

OpenStudy (anonymous):

17*17=289

OpenStudy (anonymous):

17^2 = 289

OpenStudy (anonymous):

\[(f(4))^{2}=((4)^{2}+1)^{2}= (16+1)^{2} = (17)^{2}= 289\] Hope this helps!